# Geometry-Informed Neural Operator Transformer

### v4

@@ 4 Conclusions
Extensive experiments on diverse datasets have demonstrated GINOT’s excellent accuracy and strong generalization capabilities across various geometries. These results establish GINOT as a powerful surrogate model for computational physics involving complex domains, providing significant computational speedups while maintaining high predictive accuracy. Future work will focus on extending GINOT to tackle multi-physics problems, further enhancing its utility in scientific computing and engineering applications.
@@ 5 Data availability
The dataset is available on Zenodo [36] and the trained models are available at the GitHub repository https://github.com/QibangLiu/GINOT
@@ 7 Acknowledgements
The authors would like to thank the National Center for Supercomputing Applications (NCSA) at the University of Illinois, and particularly its Research Computing Directorate, Industry Program, and Center for Artificial Intelligence Innovation (CAII) for support and hardware resources. This research used both the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois, and the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta and DeltaAI are joint efforts of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.
@@ 8 Author contributions statement
Q. Liu conceptualized the study, developed the methodology, implemented the models, prepared the datasets, conducted the analysis, and drafted, reviewed, and edited the manuscript. W. Zhong contributed to running the baseline models (GNO and GI-FNO), manuscript review and editing. H. Meidani contributed to conceptualization, results discussion, and manuscript review and editing. D. Abueidda contributed to results discussion and manuscript review and editing. S. Koric prepared the bracket lug dataset, contributed to conceptualization, results discussion, and reviewed and edited the manuscript. P. Geubelle contributed to the supervision, methodology, results discussion, and manuscript review and editing.

### v5

@@ 4 Conclusions
Extensive experiments on diverse datasets have demonstrated GINOT’s good accuracy and generalization capabilities across various geometries. These results establish GINOT as a powerful surrogate model for computational physics involving complex domains, providing significant computational speedups while maintaining high predictive accuracy. Future work will focus on extending GINOT to tackle multi-physics problems, further enhancing its utility in scientific computing and engineering applications.
@@ 5 Data availability
The dataset is available on Zenodo [liu2025geometry] and the trained models are available at the GitHub repository https://github.com/QibangLiu/GINOT
@@ 7 Acknowledgements
The authors would like to thank the National Center for Supercomputing Applications (NCSA) at the University of Illinois, and particularly its Research Computing Directorate, Industry Program, and Center for Artificial Intelligence Innovation (CAII) for support and hardware resources. This research used both the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois, and the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta and DeltaAI are joint efforts of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications. This work used Delta/DeltaAI at NCSA through allocations MAT240113 and CHE230009 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@ 8 Author contributions statement
Q. Liu conceptualized the study, developed the methodology, implemented the models, prepared the datasets, conducted the analysis, and drafted, reviewed, and edited the manuscript. W. Zhong contributed to running the baseline models, manuscript review and editing. H. Meidani contributed to conceptualization, results discussion, and manuscript review and editing. D. Abueidda contributed to results discussion and manuscript review and editing. S. Koric prepared the bracket lug dataset, contributed to conceptualization, results discussion, and reviewed and edited the manuscript. P. Geubelle contributed to the supervision, methodology, results discussion, and manuscript review and editing.